I need help with this please. Graph the parabola whose equation is given and answer the questions.
y=-x^2-4x-3
I am stuck on finding the x intercepts and y intercepts for this. I have gotten that this opens downward and the vertex is (-2,1).
Thanks!

Answer by stanbon(75887)

Graph the parabola whose equation is given and answer the questions.
y=-x^2-4x-3
I am stuck on finding the x intercepts and y intercepts for this. I have gotten that this opens downward and the vertex is (-2,1).
-----
y = -x^2 -4x -3
---
x-intercept ??
Let y = 0 ; then solve for "x";
-x^2 -4x -3 = 0
----
Divide thru by -1 to get:
x^2 + 4x + 3 = 0
-----
Factor:
(x+3)(x+1) = 0
---
Solve for "x":
x = -3 or x = -1
----
The x-intercepts are (-3,0) and (-1,0)
======
y - intercept ??
Let x = 0 ; then solve for "y":
y = -0^2 -4*0 - 3
y = -3
---
The y-intercept is (0,-3)
==============================
Your vertex work is correct.
==============================
Cheers,
Stan H.
--------------